User name jpchen
Log entry time 09:40:28 on October 27,1998
Entry number 3820
keyword=target pol measurements
8:47 turned off transverse laser (we completed transverse data taking for p=2.58 GeV).
8:50-9:00 moved into pickup position, rotated field from 270 to 0 degree.
9:05 NMR measurement without laser on:
Peak=11.94 mv, pol.=32.27/31.07%.
9:12-16 turned on longitudinal lasers 1, 2, 3.
9:20 NMR measurement with lasers on:
Peak=12.93 mv, pol=34.94/33.65%.
9:30 Moved to EMPTY position for current calibration.